OpenStudy (riley):

Calculate dy/dx. Simplify your answer. y = x^3(8-x^2)

OpenStudy (riley):

I missed one day of class this week when they went over this lesson . . . so now I'm kind of behind . . . and don't completely understand the dy/dx concept yet. >_<

OpenStudy (anonymous):

first expand = 8x^3 - x^5 dy/dx = 24x^2 - 5x^4

OpenStudy (riley):

oh I just multiply through, and then use the power rule?

OpenStudy (anonymous):

use the stabdrad form: if y = ax^n then dy/dx = anx^(n-1)

OpenStudy (anonymous):

yes

OpenStudy (riley):

Ohhhhh that makes perfect sense lol. Thank you very much!!

OpenStudy (anonymous):

stabdrad = standard!

OpenStudy (anonymous):

no probs